OpenStudy (anonymous):

Solve the problem. Make sure your answer is in simplest form and improper fractions are expressed as mixed numbers.

OpenStudy (anonymous):

\[\frac{ 1 }{ 5 } * \frac{ 6 }{ 7 } = ?\]

OpenStudy (anonymous):

@jeffyblood

OpenStudy (anonymous):

I got 6/35

OpenStudy (anonymous):

that is one of my possible answers

OpenStudy (anonymous):

I have 4 qustions left but 3 more need to correct

OpenStudy (anonymous):

what are they...

OpenStudy (anonymous):

all preety much * with fractions

OpenStudy (anonymous):

Alright give me the questions

OpenStudy (anonymous):

mk

OpenStudy (anonymous):

Which are products for this problem, including products that are not simplified? Choose all answers that are correct.

OpenStudy (anonymous):

here

OpenStudy (anonymous):

\[2\frac{ 1 }{ 3 } * \frac{ 3 }{ 4 }\]

OpenStudy (anonymous):

hold on...

OpenStudy (anonymous):

let me give you the answers

OpenStudy (anonymous):

\[1\frac{ 3 }{ 4 }\]

OpenStudy (anonymous):

\[2\frac{ 3 }{ 12 }\]

OpenStudy (anonymous):

\[\frac{ 7 }{ 4 }\]

OpenStudy (anonymous):

\[2\frac{ 1 }{ 4 }\]

OpenStudy (anonymous):

ok thats it

OpenStudy (anonymous):

1 3/4 and 7/4

OpenStudy (anonymous):

Choose all answers that are correct. means just 2 answers

OpenStudy (anonymous):

here is next

OpenStudy (anonymous):

Solve the problem. Make sure your answer is in simplest form.

OpenStudy (anonymous):

\[\frac{ 3 }{ 16 } * \frac{ 4 }{ 21 }\]

OpenStudy (anonymous):

:-I

OpenStudy (anonymous):

1/28 is the answer. my bad I took a while

OpenStudy (anonymous):

ok me pc lagy so I thoght it was that so ya thx 2 more

OpenStudy (anonymous):

Solve the problem. Make sure your answer is in simplest form and improper fractions are expressed as mixed numbers.

OpenStudy (anonymous):

where they right?

OpenStudy (anonymous):

\[2\frac{ 2 }{ 5 } * 2\frac{ 7 }{ 9 }\]

OpenStudy (anonymous):

6 2/3

OpenStudy (anonymous):

ok thx

OpenStudy (anonymous):

last one

OpenStudy (anonymous):

Solve the problem. Make sure your answer is in simplest form and improper fractions are expressed as mixed numbers.

OpenStudy (anonymous):

A packer fills \[8\frac{ 1 }{ 4 }\] boxes in 1 hour. How many boxes can the packer fill in\[7\frac{ 1 }{ 3 }\] hours?

OpenStudy (anonymous):

@Unanimoose

OpenStudy (anonymous):

My bad I went to lunch